# Scaling up to Multivariate Rational Function Reconstruction

### v1

@@ 1 Introduction
The current standard approach is to derive integration-by-parts identities [1] for a set of seed integrals with fixed powers of propagators and irreducible scalar products and solve the resulting system of linear relations via Gauss elimination [2]. The result is a subset of the seed integrals, each expressed in terms of a linear combination of basis integrals. The coefficients are rational functions of kinematic invariants and the space-time dimension.
It is often advantageous to insert numerical values for the dimension and the invariants and solve the system over a finite field [3]. This strategy was initially used to quickly eliminate redundant relations [4]. However, solving the system with sufficiently many different probes, i.e. different numeric values for the variables, it is possible to reconstruct the full result [5, 6]. In this way, one avoids large intermediate expressions and can restrict the reconstruction to those coefficients that are actually needed for the final result, for example a scattering amplitude.
This general strategy has been optimised along several directions. Right from the start, one can attempt to find combinations of relations that lead to better behaved systems or even an explicit recursive solution to the reduction problem [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Since in a finite-field reduction the very same system of linear equations has to be solved many times, it can be worthwhile to record the solution steps once, optimise the recorded sequence, and replay it to rapidly obtain further probes [19]. Other efforts target the reconstruction itself. The complexity of the rational function coefficients depends on the chosen basis integrals, and a judicious choice results in a factorisation between the dimension and the kinematic invariants [20, 21]. Identifying common factors in the coefficients can further reduce the number of probes required for the reconstruction [22, 23, 24, 25]. Finally, the number of required probes depends on the chosen reconstruction method, and various algorithms with different strengths have been explored [26, 27, 28].
@@ 2 Univariate Rational Function Reconstruction
Note that the reconstruction is optimal if two conditions are fulfilled. First, the degrees of the numerator and the denominator in $f$ should be equal or the degree of the denominator should be larger by one. Second, the numerator and denominator polynomials should be perfectly dense, with all coefficients non-zero. In this case the number of required probes is equal to the number of unknown coefficients plus the chosen number of probes used for confirmation. Empirically, the rational functions encountered in integration-by-parts reduction without any kinematic invariants are close to ideal with a typical overhead of about 10% in the number of required probes.
@@ 3 Scaling up to Multiple Variables
where the sums run over all powers $0\leq p_{1}\leq P_{1},\dots,0\leq p_{n}\leq P_{n}$ in the numerator and $0\leq q_{1}\leq Q_{1},\dots,0\leq q_{n}\leq Q_{n}$ in the denominator. The degrees $P_{1},\dots,P_{n}$ and $Q_{1},\dots,Q_{n}$ are a priori unknown. To reconstruct a pair $(P_{i},Q_{i})$ of degrees we can set all other variables to some fixed value, i.e. $x_{j}=t_{j}$ for all $j\neq i$, and perform a univariate rational function reconstruction in the remaining free variable $x_{i}$.
@@
The method described so far can suffer from accidental cancellations between numerator and denominator. For example, for $f(x_{1},x_{2})=\frac{x_{2}}{x_{1}}$ we would obtain the auxillary function $g(x)=x$, which would lead us to believe the original function was $f(x_{1},x_{2})=x_{1}$. To prevent this, we additionally shift the rescaled argument by a randomly chosen number.
@@ 4 Application to the Reduction to Basis Integrals
Let us now assess the efficiency of the algorithm presented in section 3 in practical applications. For brevity, we will refer to the new method as “scaling” reconstruction. We compare it to an implementation of an algorithm proposed by Cuyt and Lee [31]. This algorithm is described in detail in [6, 26]. We briefly recall the main steps. First, the variables are rescaled with a common factor $t$ and shifted, leading to $(x_{1},\dots,x_{n})$ = $(ty_{1}+s_{1},\dots,ty_{n}+s_{n})$ with $y_{n}=1$. One then performs a univariate rational reconstruction in $t$. Starting from the highest powers, the coefficients of $t^{i}$ are reconstructed as polynomials in $y_{1},\dots,y_{n}$ and transformed back to the original variables. Here, we are mainly interested in dense rational functions and therefore use nested Newton interpolation for the polynomial reconstruction. We call this “dense homogeneous” reconstruction, adopting terminology from [28].
[…]
@@ 4.1 Massive Four-Loop Propagator
with the ellipsis indicating a linear combination of further basis integrals with less complex coefficients. $q(z,d)$ is a dense rational function in $z$ and the space-time dimension $d$. The numerator degrees in $z$ and $d$ are $P_{z}=P_{d}=81$ and the denominator degrees are $Q_{z}=80$ and $Q_{d}=78$. The number of probes per characteristic required by the dense homogeneous and the scaling algorithm are shown in table 1 together with the number of probes a hypothetical optimal algorithm that can determine one unknown coefficient per probe would use. For a full reconstruction, probes are needed in seven different prime fields, with two probes in an eighth field used for confirmation. The scaling method requires $164$ additional probes in the first prime field to determine the degrees in $z$.
Algorithm Number of Probes
Dense Homogeneous 25 7082570825\,70825 708
Scaling 13 4301343013\,43013 430
Optimal 12 7211272112\,72112 721
Table 1: Number of probes per characteristic required to reconstruct a specific coefficient in the differential equation for a four-loop propagator.
We see that the scaling algorithm performs close to optimal, with an overhead of about $6\%$ additional probes. It requires roughly $48\%$ fewer probes than dense homogeneous reconstruction.
@@ 4.2 Diphoton Plus Jet Production at Two Loops
Next, let us consider the two-loop amplitude for diphoton plus jet production, taken from [30]. Specifically, we choose the parity-even contribution with a left-handed quark and a gluon in the initial state, a negative gluon helicity, opposite-sign photon helicities, and no closed fermion loops. Denoting the number of colours by $N_{C}$, we select the largest rational coefficient by Mathematica’s ByteCount in the part proportional to $N_{C}^{0}$. We write this coefficient as a rational function in $x_{23},x_{34},x_{45},x_{51}$, where $x_{ij}=\frac{s_{ij}}{s_{12}}$, and
@@
are Mandelstam invariants. In this example, the full coefficient can be reconstructed using a single prime field. The number of required probes is shown in table 2.
Algorithm # Probes
Homogeneous 248 688248688248\,688248 688
Scaling 169 132169132169\,132169 132
Optimal 30 4903049030\,49030 490
@@
The scaling algorithm introduced in section 3 performs notably better than the homogeneous one, requiring approximately $32\%$ fewer probes. Still, both algorithms are far from optimal for this scenario and algorithms for sparse reconstruction [26, 27, 36, 37, 38] will likely perform much better.
@@ 5 Conclusion
I have presented an algorithm for the reconstruction of dense multivariate rational functions. Multiple variables are mapped onto a single variable, using scaling powers and shifts chosen such that the mapping can be reversed. In this way, the problem is reduced to well-known univariate rational reconstruction.
The algorithm is tested on two examples taken from complex reductions to basis integrals, a massive four-loop propagator and and two-loop five-point amplitude. For the dense rational function encountered in the four-loop problem, the required number of probes exceeds the number of unknown coefficients by only about $6\%$. In the sparse two-loop example, the overhead is much larger and the number of probes exceeds the number of non-vanishing coefficients by a factor of more than $5$. Still, in both cases the scaling algorithm compares favourably to dense homogeneous reconstruction [31]. A promising avenue for future research would be to combine the univariate mapping with sparse rational reconstruction in a single variable [39].
@@ Acknowledgements
I thank P. Marquard for enlightening discussions and helpful comments on the manuscript.

### v2

@@ 1 Introduction
The current standard approach is to derive integration-by-parts identities [1, 2] for a set of seed integrals with fixed powers of propagators and irreducible scalar products and solve the resulting system of linear relations via Gauss elimination [3]. The result is a subset of the seed integrals, each expressed in terms of a linear combination of basis integrals. The coefficients are rational functions of kinematic invariants and the space-time dimension.
It is often advantageous to insert numerical values for the dimension and the invariants and solve the system over a finite field [4]. This strategy was initially used to quickly eliminate redundant relations [5]. However, solving the system with sufficiently many different probes, i.e. different numeric values for the variables, it is possible to reconstruct the full result [6, 7]. In this way, one avoids large intermediate expressions and can restrict the reconstruction to those coefficients that are actually needed for the final result, for example a scattering amplitude. Further advantages include ease of parallelisation, lower memory usage, and the possibility to optimise the system further after a computationally cheap pilot run.
This general strategy has been further developed along several directions. Right from the start, one can attempt to find combinations of relations that lead to better behaved systems or even an explicit recursive solution to the reduction problem [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Since in a finite-field reduction the very same system of linear equations has to be solved many times, it can be worthwhile to record the solution steps once, optimise the recorded sequence, and replay it to rapidly obtain further probes [20]. Other efforts target the reconstruction itself. The complexity of the rational function coefficients depends on the chosen basis integrals, and a judicious choice results in a factorisation between the dimension and the kinematic invariants [21, 22]. Identifying common factors in the coefficients can further reduce the number of probes required for the reconstruction [23, 24, 25, 26]. Finally, the number of required probes depends on the chosen reconstruction method, and various algorithms with different strengths have been explored [27, 28, 29, 30].
@@ 2 Univariate Rational Function Reconstruction
Note that the reconstruction is optimal if two conditions are fulfilled. First, the degrees of the numerator and the denominator in $f$ should be equal or the degree of the numerator should be larger by one. Second, the numerator and denominator polynomials should be perfectly dense, with all coefficients non-zero. In this case the number of required probes is equal to the number of unknown coefficients plus the chosen number of probes used for confirmation. Empirically, the rational functions encountered in integration-by-parts reduction without any kinematic invariants are close to ideal with a typical overhead of about 10% in the number of required probes.
@@ 3 Scaling up to Multiple Variables
where the sums run over all powers $0\leq p_{1}\leq P_{1},\dots,0\leq p_{n}\leq P_{n}$ in the numerator and $0\leq q_{1}\leq Q_{1},\dots,0\leq q_{n}\leq Q_{n}$ in the denominator. The degrees $P_{1},\dots,P_{n}$ and $Q_{1},\dots,Q_{n}$ are a priori unknown. To reconstruct a pair $(P_{i},Q_{i})$ of degrees we can set all other variables to some fixed value, i.e. $x_{j}=t_{j}$ for all $j\neq i$, and perform a univariate rational function reconstruction in the remaining free variable $x_{i}$ [7, 27, 33].
@@
The method described so far can suffer from accidental cancellations between numerator and denominator. For example, for $f(x_{1},x_{2})=\frac{x_{2}}{x_{1}}$ we would obtain the auxillary function $g(x)=x$, which would lead us to believe the original function was $f(x_{1},x_{2})=x_{1}$. To prevent this, we additionally shift the rescaled argument by a randomly chosen number. [^1] This random shift is also used with a slightly different purpose in the reconstruction algorithm by Cuyt and Lee [7, 27, 33]. There, the goal is to ensure a unique structure and uniform coefficient normalisation of the reconstructed function. Spurious cancellations have to involve two or more different variables. We therefore expect to avoid them by having at least one shifted variable in each possible combination, i.e. we shift each variable except one.
@@
One of the main advantages of numerical reduction is ease of parallelisation. This requires that subsequent probes can be chosen without having to wait for the outcome of feeding earlier probes into the reconstruction algorithm. In this respect, algorithms for the reconstruction of dense rational functions tend to perform better than methods aimed at sparse rational functions with many vanishing coefficients, as for the latter the probe selection typically has to be adjusted dynamically. The presented algorithm mostly decouples the seed choice from the reconstruction progress. For $n$ variables, the selection strategy has to be updated $n$ times — after determining the powers of each of the first n1n-1 variables and once more to use the final rescaled arguments.
@@ 4 Application to the Reduction to Basis Integrals
Let us now assess the efficiency of the algorithm presented in section 3 in practical applications. For brevity, we will refer to the new method as “scaling” reconstruction. We compare it to an algorithm proposed by Cuyt and Lee [33]. This algorithm is described in detail in [7, 27]. We briefly recall the main steps. First, the variables are rescaled with a common factor $t$ and shifted, leading to $(x_{1},\dots,x_{n})$ = $(ty_{1}+s_{1},\dots,ty_{n}+s_{n})$ with $y_{n}=1$. One then performs a univariate rational reconstruction in $t$. Starting from the highest powers, the coefficients of $t^{i}$ are reconstructed as polynomials in $y_{1},\dots,y_{n}$ and transformed back to the original variables. For the comparison we use state-of-the-art implementations in the public codes FireFly [27, 29, 34] and FiniteFlow [28, 35]. The comparison code and the example rational functions in computer-readable form are available from https://github.com/a-maier/scaling-rec.
For the scaling algorithm, the reconstruction is first performed over a number of prime fields $\mathbb{Z}_{P}$, using arithmetic algorithms taken from NTL [36]. We start with P=1 152 921 504 606 846 883P=\$1\,152\,921\,504\,606\,846\,883\$ and move to the next smaller prime numbers as needed. The resulting coefficients are lifted to a higher characteristic with the Chinese remainder theorem, specifically Bézout’s identity. The actual rational coefficients are then reconstructed from the finite-field integers via an algorithm by Wang [37]. Again, details are given in [7, 27].
In principle, the reconstruction can be simplified tremendously exploiting the structure of the result [23, 24, 25]. Both FireFly and FiniteFlow can factorise the numerator and denominator to a certain degree. FiniteFlow determines the minimal degree in each variable to automatically factor out common monomials. FireFly optionally performs univariate factorisation to identify any factors depending on a single variable. Since we are mainly interested in assessing the underlying reconstruction algorithms, we disable FireFly’s factorisation in the following comparisons. As there is no option to switch off factorisation in FiniteFlow, we additionally compare the reconstruction after removing all monomial factors from the function to be reconstructed.
@@ 4.1 Massive Four-Loop Propagator
with the ellipsis indicating a linear combination of further basis integrals with less complex coefficients. $q(z,d)$ is a rational function in $z$ and the space-time dimension $d$, where the numerator degrees in $z$ and $d$ are $P_{z}=P_{d}=81$ and the denominator degrees are $Q_{z}=80$ and $Q_{d}=78$. The numbers of probes required for reconstructing the function over the first characteristic are shown in table 1 for the scaling algorithm and FireFly together with a hypothetical optimal algorithm that can determine one unknown coefficient per probe.
For a full rational function reconstruction probes are needed in several additional prime fields. Since different implementations vary vastly in the amount of reused information we refrain from a quantitative comparison.
Method Number of Probes
This work   13 59413\,594
FireFly   16 37316\,373
Optimal   12 72112\,721
Table 1: Number of probes required to reconstruct a specific coefficient in the differential equation for a four-loop propagator over the first characteristic.
Comparing the “Optimal” entry of table 1 to the total number of 13 12313\,123 monomials in the ansatz given by equation (8) we observe that $q(z,d)$ is dense in the sense that about $97\%$ of the coefficients in the ansatz are non-zero. We see that the scaling algorithm performs close to optimal, with an overhead of about $7\%$ additional probes. In comparison, FireFly requires approximately $29\%$ more probes.
The denominator of the reconstructed function contains an overall factor of $d^{1}z^{3}$. The improvement gained by identifying and removing this factor is illustrated in table 2, where we now also include FiniteFlow. The FiniteFlow entry does not include a few hundred probes used to determine the overall degree of the rational function and the degrees with respect to the individual variables from Thiele interpolation, c.f. section 2 [40]. The other entries count the total number of function evaluations.
Method Number of Probes
This work   13 59413\,594
FireFly   16 02016\,020
FiniteFlow $\gtrsim{}$18\,205$\$
Optimal   12 72112\,721
Table 2: Number of probes required to reconstruct a specific coefficient in the differential equation for a four-loop propagator over the first characteristic after removing an overall monomial factor. The FiniteFlow entry does not include a few hundred probes used for degree determinations.
For the algorithm presented in section 3, the scaling powers in equation (10) are completely determined by the numerator in the present example. Thus, removing factors from the denominator does not affect the number of probes needed. However, we do observe a slight reduction in the number of required evaluations with FireFly, reducing the overhead to $26\%$ compared to the optimum. The number of evaluations needed with FiniteFlow exceeds the number of non-vanishing coefficients to be determined by about $43\%$.
@@ 4.2 Diphoton Plus Jet Production at Two Loops
Next, let us consider the two-loop amplitude for diphoton plus jet production, taken from [32]. Specifically, we choose the parity-even contribution with a left-handed quark and a gluon in the initial state, a negative gluon helicity, opposite-sign photon helicities, and no closed fermion loops. Denoting the quark helicity by $\lambda_{q}$, the number of active flavours by $n_{f}$, the number of colours by $N_{C}$, and the parity transformation operator by $\hat{P}$, the reduction has the structure
[…]
where $c_{-2},c_{0},c_{2}$ are linear combinations of pentagon functions [41] with rational coefficients. From $c_{0}$ we select the largest of these coefficients by Mathematica’s ByteCount. We write this coefficient as a rational function in $x_{23},x_{34},x_{45},x_{51}$, where $x_{ij}=\frac{s_{ij}}{s_{12}}$, and
@@
are Mandelstam invariants. After determining the numerator and denominator degrees our ansatz according to equation (8) contains 136 934136\,934 unknown coefficients. However, the actual rational function is much sparser than in the example in section 4.1 and only approximately $22\%$ of these coefficients are non-zero. In this example, the full coefficient can be reconstructed using a single prime field. We collect the number of required probes in table 3.
Method Number of Probes
This work   169 132169\,132
FireFly   163 094163\,094
Optimal   30 49030\,490
@@
The scaling algorithm introduced in section 3 performs slightly worse than FireFly’s reconstruction. Both algorithms are far from optimal for this scenario, requiring more than five probes for each unknown coefficient.
The number of required reconstruction probes after removing an overall monomial factor $x_{23}^{2}x_{34}^{2}x_{45}^{2}x_{51}^{2}$ is shown in table 4. As in section 4.1, we see no improvement for the implementation of the algorithm presented in this work. In contrast, FireFly needs approximately $20\%$ fewer function evaluations than before. Most strikingly, FiniteFlow is much closer to optimal than both FireFly and the scaling reconstruction implementation, especially when using the FFPolyVandermonde alternative polynomial reconstruction method. Even when enabling its identification of univariate factors, FireFly still requires 87 48587\,485 probes, substantially more than FiniteFlow. This difference between FiniteFlow and FireFly is unexpected and deserves closer inspection. However, since the focus of the present work is on the scaling algorithm and dense reconstruction, we leave further investigation to future work.
Method Number of Probes
This work   169 132169\,132
FireFly   129 894129\,894
FiniteFlow $\gtrsim$49\,216$\$
FiniteFlow with FFPolyVandermonde $\gtrsim$47\,381$\$
Optimal   30 49030\,490
Table 4: Number of required probes after removing the overall monomial prefactor. The FiniteFlow entries do not include a few hundred probes used for degree determinations.
@@ 5 Conclusion
I have presented an algorithm for the reconstruction of dense multivariate rational functions. Multiple variables are mapped onto a single variable, using scaling powers and shifts chosen such that the mapping can be inverted. In this way, the problem is reduced to well-known univariate rational reconstruction.
The algorithm is tested on two examples taken from complex reductions to basis integrals, a massive four-loop propagator and a two-loop five-point amplitude. For the dense rational function encountered in the four-loop problem, the required number of probes exceeds the number of unknown coefficients by only about $7\%$. This compares favourably with the current state-of-the-art programs FireFly [27, 29] and FiniteFlow [28].
In the sparse two-loop example, the number of probes needed is about $4\%$ above the FireFly result when disabling factorisation. However, a comparison to FiniteFlow reveals that in this case there is substantial room for improvements for both FireFly and the presented algorithm. A further promising avenue for future research would be to combine the univariate mapping with sparse rational reconstruction in a single variable, see e.g. [42].
@@ Acknowledgements
I thank P. Marquard for enlightening discussions and helpful comments on the manuscript. I further thank F. Lange and T. Peraro for communication on FireFly and FiniteFlow and for contributing code to the comparison in section 4.